OpenStudy (anonymous):

Karen bakes cookies. She has observed that the number of cookies that turn out well in each batch varies with the size of the batch and is given by the function f(x) = 2x – 23, where x is the size of the batch (number of cookies in each batch). She never bakes fewer than 12 cookies in a batch. If the size of five batches is {20, 25, 26, 27, 34}, how many cookies turn out well in each batch? a){17, 27, 29, 31, 45} b){18, 24, 26, 30, 46} c){14, 23, 25, 31, 42} d){20, 24, 25, 26, 44} please help !

OpenStudy (anonymous):

you're gonna have to plug in the size of the five batches they give you in five separate equations. f(20) = 2(20) - 23 and f(25) = 2(25) - 23 and so on

OpenStudy (anonymous):

so its a ?

OpenStudy (anonymous):

yes its a